OpenStudy (anonymous):

what is the slope of the line given by the equation ? y=12x -23

hartnn (hartnn):

compare with y=mx+c m is the slope.

OpenStudy (mathstudent55):

An equation in the form y = mx + b is said to be in the slope-intercept form. Your equation is in this form since you can match exactly y = 12x - 23 to y = mx + b. In the slope-intercept form, y = mx + b, m is the slope, and b is the y-intercept, or where the line crosses the y-axis. In your case, m = 12, so the slope is 12. The y-intercept is -23, which means that the line crosses the y-axis at y = -23, which is the point (0, -23).